OpenStudy (anonymous):

differentiate the problem g(x)=4/3x^3

OpenStudy (john_es):

You mean calculate the derivative? \[g(x)=3\cdot\frac{4}{3}x^{3-1}=4x^2\]

OpenStudy (john_es):

In g(x) I mean, \[g'(x)\]

OpenStudy (anonymous):

Thats not the answer given in the book its -4x^-4 probably the reason im confused

OpenStudy (john_es):

Not, may be you mean differentiate the function, \[\frac{4}{3x^3}\]

OpenStudy (john_es):

Then the derivative is, \[\left(\frac{4}{3}x^{-3}\right)'=\frac{4}{3}(-3)x^{-3-1}=-4x^{-4}\]

OpenStudy (anonymous):

Thank you so here is another one f(x)= x/7+ 7/x

OpenStudy (john_es):

\[\left(\frac{x}{7}+\frac{7}{x}\right)'=\left(\frac{x}{7}\right)'+\left(\frac{7}{x}\right)'=\frac{1}{7}+(7x^{-1})'=\frac{1}{7}-7x^{-2}=\frac{1}{7}-\frac{7}{x^2}\]

OpenStudy (anonymous):

G(x)=1/3Г8x^2

OpenStudy (john_es):

There is a symbol I don't understand, between the 3 and the 8.

OpenStudy (anonymous):

Closest could get to a square root

OpenStudy (john_es):

Ok so it is, \[G(x)=\frac{1}{3\sqrt{8x^2}}\]?

OpenStudy (john_es):

Or it is a cubic root? \[G(x)=\frac{1}{\sqrt[3]{8x^2}}\]

OpenStudy (anonymous):

Sorry cube root

OpenStudy (john_es):

Ok, then \[G(x)=(8x^2)^{-1/3}\Rightarrow G(x)=\frac{1}{2}x^{-2/3}\Rightarrow \\ \Rightarrow G'(x)=\frac{1}{2}\frac{-2}{3}x^{-2/3-1}=-\frac{1}{3}x^{-5/3}\]

OpenStudy (anonymous):

X^3 cubic root x

OpenStudy (john_es):

You mean, \[x^3\sqrt[3]{x}\]?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Sorry to be confusing no computer

OpenStudy (john_es):

Ok, then product rule, \[(x^3\sqrt[3]{x})'=(x^3)'\sqrt[3]{x}+x^3(\sqrt[3]{x})'=3x^2\sqrt[3]x+x^3\frac{1}{3}x^{-2/3}=3x^{7/3}+\frac{1}{3}x^{7/3}=\frac{10}{3}x^{7/3}\]

OpenStudy (anonymous):

\[\eqalign{ &\frac{d}{dx}\frac{4x^3}{3}\\ =&\frac{4}{3}\frac{d}{dx}x^3 \\ =&\frac{4}{3}(3x^2) \\ =&4x^2 }\]

OpenStudy (anonymous):

Can you explain that one ?

OpenStudy (john_es):

Another way without product rule, \[x^3\sqrt[3]{x}=x^{10/3}\Rightarrow (x^{10/3})'=\frac{10}{3}x^{10/3-1}=\frac{10}{3}x^{7/3}\]

OpenStudy (john_es):

Remember, \[x^3\sqrt[3]x=x^3\cdot x^{1/3}=x^{3+1/3}=x^{10/3}\]

OpenStudy (anonymous):

Ok makes sense now x^3(3x)^2

OpenStudy (john_es):

Simplifying \[x^3(3x)^2=x^39x^2=9x^5\Rightarrow(9x^5)'=45x^4\]

OpenStudy (anonymous):

Ok sorry been while for me f(q)=3q^2+4q-2over q

OpenStudy (anonymous):

?

OpenStudy (john_es):

You mean, \[\frac{3q^2+4q-2}{q}\]?

OpenStudy (anonymous):

Yes

OpenStudy (john_es):

Then simplify first, \[3q+4-2/q\]Then derive, \[(3q+4-2/q)'=3+2/q^2\]

OpenStudy (anonymous):

X^2+x^3over x ^2